OpenStudy (anonymous):

Solve system by elimination: -2x+4y=20 5x+8y=2

OpenStudy (lgbasallote):

multiply first eq by -2 4x -8y = -40 5x + 8y = 2 9x = -38 x = -38/9 to get y...multiply first eq by 5 the second by 2 -10x + 20 y = 100 10x + 16y = 4 36y = 104 y = 104/36 y = 26/9